I am very confused as to how to Solve each system by the additional method. Can someone help me understand this and solve it please?
17.) x + y = 48
12x + 14y = 628

Answer by arunpaul(104)

x + y = 48 ------eq1
12x + 14y = 628 ------eq2
12*eq1 -eq2
so
12x +12 y = 576
12x + 14y = 628
------------------(subtracting we have)
-2y =(576-628)=-52
y = 52/2=26 (Ans)